Hi I'm an older student and I have a problem I don't understand. How would you solve polynomials by factoring? This is the problem: 2x^4=16x^2. Thank You
Answer by AnlytcPhil(1116)

Hi I'm an older student and I have a problem I don't
understand. How would you solve polynomials by factoring?
This is the problem: 2x^4=16x^2. Thank You

                  2x4 = 16x2 
First observe that both sides can be divided by 2. So we
can make it simpler by dividing both sides by 2
                   x4 = 8x2
Get zero on the right side:
            x4 - 8x2 = 0
Factor out x2 on the left side:
           x2(x2 - 8) = 0
Set each of these factor = 0 separately and solve:
Setting first factor = 0
                    x2 = 0
                    x = 0
Setting second factor = 0
               x2 - 8 = 0
                   x2 = 8
Take square roots of both sides

` ` ` ` ` ` ` ` ` ` ` ` ` _
` ` ` ` ` ` ` ` ` ` x = ±Ö8
Simplify radical:` ` ` ` `___
` ` ` ` ` ` ` ` ` ` x = ±Ö4·2
` ` ` ` ` ` ` ` ` ` ` ` ` _ `_
` ` ` ` ` ` ` ` ` ` x = ±ÖÖ2
` ` ` ` ` ` ` ` ` ` ` ` ` `_
` ` ` ` ` ` ` ` ` ` x = ±2Ö2
`
Edwin J